OpenStudy (anonymous):

You are playing a game with a friend where you flip a coin and if it comes up heads you give her 1 dollar and if it comes up tails she gives you one dollar. What would be the variance of your earnings?

OpenStudy (anonymous):

Let x = number of times it comes up tails. then x follows a binomial distribution (n, p) where n is the nuber of times you flip and p = P(T) on a flip. your earnings will be x - (n-x) = 2x-n Variance of binomial r.v. is n*p*(1-p). variance of a constant is 0. you need the VAR(2x-n) = VAR(2x)-VAR(n) = VAR(2x) = 4*VAR(x) = 4*n*p*(1-p)